If Qs = -20 + 10p, and Qd = 400 - 20p, what is the equilibrium price and quantity?


1
Expert's answer
2021-06-13T17:39:09-0400

If Qs = -20 + 10p, and Qd = 400 - 20p, then in equilibrium Qd = Qs, so:

"-20 + 10p = 400 - 20p,"

30p = 420,

p = 14,

"Q = -20 + 10\u00d714 = 120" units.